# GT2-GS: Geometry-aware Texture Transfer for Gaussian Splatting

### v3

@@ Introduction
3D style transfer (zhang2022arf; zhang2024stylizedgs) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information. With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has increased significantly. 3D stylization techniques offer a promising solution by accelerating the creation of 3D content, particularly in complex 3D scene environments.
Recently, the emergence of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl3Dgaussians) has significantly advanced the field of 3D stylization. NeRF-based stylization methods (zhang2022arf; zhang2023ref; nguyen2022snerf), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry. In contrast, 3DGS-based approaches (galerne2025sgsst; liu2025abc; liu2024stylegaussian) offer benefits such as explicit editability and real-time rendering. 3D stylization methods can be broadly categorized into feed-forward and optimization-based approaches, depending on whether they support zero-shot style transfer. While feed-forward methods (liu2023stylerf; huang2022stylizednerf; liu2024stylegaussian) enable zero-shot stylization, they often exhibit lower rendering quality. In contrast, optimization-based methods (zhang2022arf; liu2025abc; zhang2024stylizedgs; galerne2025sgsst) typically produce higher-fidelity stylization results. Despite these differences, both types of methods commonly define optimization objectives within the VGG feature space during training.
@@
Early texture transfer algorithms (ashikhmin2003fast; lee2010directional; hertzmann2023image; efros2023image) were based on non-parametric texture synthesis. These methods preserve the structural information of the target image to achieve texture transfer effects. Gatys et al. (gatys2016image) pioneered the use of convolutional neural networks (CNNs) (simonyan2014very) for image style transfer, achieving impressive results. Subsequently, a variety of deep learning-based style transfer methods (chen2016fast; li2016combining; huang2017arbitrary; huo2021manifold) have been proposed. These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer. Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (vaswani2017attention), some deep learning-based texture transfer methods (wang2022texture; chen2022auv; pu2024dynamic) can transfer local textures from the reference image to the target image based on high-level semantic correspondence. In this work, we explore how to directly transfer texture features into 3D representations.
@@
3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency. Early research in 3D scene style transfer focused on explicit representations such as point clouds (cao2020psnet; huang2021learning), meshes (hollein2022stylemesh; yin20213dstylenet), and voxels (guo2021volumetric; klehm2014property). Zhang et al. (zhang2022arf) explored the potential of NeRF in 3D stylization tasks. By optimizing the scene representation with the proposed 3D style transfer loss, it achieved pleasing visual effects. Subsequently, a variety of NeRF-based 3D style transfer methods (huang2022stylizednerf; pang2023locally; liu2023stylerf; zhang2024coarf; jung2024geometry) have emerged. Recently, the emergence of 3DGS (kerbl3Dgaussians) has brought new possibilities for scene stylization. It features fast training speed, high rendering quality, and efficient performance. Some works (liu2024stylegaussian; zhang2024stylizedgs; mei2024regs; liu2025abc) have already explored the potential of stylization under this representation. Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (liu2024stylegaussian) and optimization-based approaches (zhang2024stylizedgs; mei2024regs; liu2025abc). Feed-forward methods enable zero-shot style transfer. For instance, StyleGaussian (liu2024stylegaussian) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (huang2017arbitrary) to achieve real-time stylization. On the other hand, optimization-based methods attain higher-quality style transfer results. Some works (zhang2024stylizedgs; liu2025abc) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization. However, existing methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images. Meanwhile, these methods overlook the fine-grained discrepancies between features and pixels. By incorporating geometric information and adaptive fine-grained control, our method is able to achieve high-quality texture transfer results.
@@
3D Gaussian Splatting (kerbl3Dgaussians) is a novel explicit 3D representation method. It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes. The parameters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, an opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$, which is represented as spherical harmonic coefficients. Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process. The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$. The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (lassner2021pulsar). Specifically, the Gaussians are grouped into tiles and sorted by depth. The color $C$ of a pixel is computed through $\alpha$-blending:
@@
By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized. Existing optimization-based 3D style transfer methods (zhang2022arf; pang2023locally; zhang2024coarf; zhang2024stylizedgs) typically adopt the nearest neighbor feature matching (NNFM) loss (zhang2022arf) as their style loss function. It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them. Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$. The same feature extractor (e.g., VGG (simonyan2014very)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively. Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
@@
Datasets. For the scene datasets, we utilize the LLFF dataset (mildenhall2019local) and the Tanks and Temples (T&T) dataset (knapitsch2017tanks), which are collected from the real world. Additionally, we use images from the ARF (zhang2022arf) style dataset and the DTD dataset (cimpoi2014describing) as reference image datasets.
Baseline. We compare our method with the state-of-the-art 3D stylization methods, including SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref), and SNeRF (nguyen2022snerf). Specifically, SGSST, ABC-GS, and StyleGaussian are based on 3DGS, and ARF, Ref-NPR, and SNeRF are based on NeRF. StyleGaussian is a feed-forward-based method, while others are optimization-based methods.
@@
We perform view-consistent color transfer (zhang2022arf) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters. For the VGG (simonyan2014very) feature extractor, we employ the conv3 block of VGG-16. When associating color parameters with texture features, we set the default number of depth groups $K$ to 4, and the rotation angle $\theta$ is sampled across the full 360 degrees. The hyperparameter for AFCM is denoted as $\{\lambda_{d},\lambda_{f},\lambda_{\Phi}\}$ = {0.8, 0.8, 0.25}. For texture transfer optimization, we set $\{\lambda_{wgt},\lambda_{c},\lambda_{tv}\}$ = {2, 0.005, 0.02}. All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@
Multi-view Consistency. We evaluate multi-view consistency (liu2024stylegaussian) using both short-term and long-term consistency metrics. The results are shown in Tab. 1. It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency. Notably, ABC-GS disables the Gaussian densification strategy during optimization. In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPB.
Content Preservation. For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance. We use SSIM (wang2004image) and CLIP-score (radford2021learning) to assess the preservation of content information. Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level. In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images. As shown in Tab. 1, our proposed method achieves significantly higher scores on both evaluation metrics. In particular, our method outperforms previous approaches in terms of the CLIP-score. It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.
@@ Acknowledgments
ADDITIONAL ANALYSIS
User Study. Given the highly subjective nature of texture transfer evaluation, it is challenging to rely solely on objective quantitative metrics for a comprehensive assessment. Therefore, we conducted a user study to evaluate the effectiveness of the proposed method in terms of texture transfer quality. We collected 25 responses from an online questionnaire distributed through social media platforms. The questionnaire consisted of 20 groups of images along with corresponding prompts. Each group included a texture image, an original content image from a random viewpoint, and a rendered image produced by one of the methods under comparison from the same viewpoint. Each rendered image was evaluated in terms of texture alignment and visual quality, with scores ranging from 1 (lowest) to 5 (highest). As shown in Tab. 3, our method consistently outperforms the baseline approaches. Specifically, by comparing the differences across various evaluation dimensions, it is evident that users showed a clear preference for our proposed method. This indicates that our approach achieves texture transfer results that better align with human visual perception.
Runtime Cost. Runtime efficiency is crucial for 3D assets. Therefore, we evaluate the runtime cost of different methods on the LLFF dataset. As shown in Tab. 4, our method demonstrates superior efficiency in terms of both memory usage and FPS. StyleGaussian incurs higher memory costs due to the use of additional VGG feature parameters during rendering. ARF, Ref-NPR, and SNeRF are NeRF-based methods, which result in slower rendering speeds.
Impact of Content Loss. As shown in Fig. 7, the content loss controls the strength of texture transfer, helping preserve the original scene content. However, an excessively large $\lambda_{c}$ hinders the scene from effectively learning the desired texture appearance.
MORE IMPLEMENTATION DETAILS
The frequency map is derived from high-frequency energy statistics based on the Discrete Cosine Transform (DCT). Specifically, according to the scale difference between the image and the feature map, the image is divided into 8×8 blocks, and DCT is performed on each block. The frequency density of a given region is defined as the sum of the absolute DCT coefficients within the corresponding block. For the forward-facing scene dataset LLFF and the 360-degree scene dataset T&T, we adopt different experimental settings. In the color transfer stage, we perform 400 and 1000 optimization steps for LLFF and T&T scenes, respectively. During the subsequent texture transfer stage, for the LLFF dataset, we set $\{\lambda_{gt},\lambda_{c},\lambda_{tv},\lambda_{reg}\}$ = {2, 0.005, 0.02, 0.1}. For the T&T dataset, we reduce parameter $\lambda_{c}$ to 0.0005. To simplify the implementation of the geometry preservation branch, our code adopts an alternating optimization strategy between the two branches.
MORE QUALITATIVE EVALUATION
Texture Orientation Control. The proposed GT Loss enables the scene to propagate the target texture information based on the prior view. Therefore, by controlling the learnable texture direction in the initial prior view, we can influence the overall appearance learned by the entire scene. This control is achieved by injecting pseudo prior angle information into the first prior view. As shown in Fig. 8, by combining the reference texture image and view direction control, we achieve high-quality texture transfer results with controllable texture orientation.
More Comparison. For more qualitative evaluation, we compare our method with SGSST (galerne2025sgsst), ABC-GS (liu2025abc), StyleGaussian (liu2024stylegaussian), ARF (zhang2022arf), Ref-NPR (zhang2023ref) and SNeRF (nguyen2022snerf). Fig. 9 and Fig. 10 present the qualitative results.
Ours SGSST ABC-GS ARF
Texture Alignment($\uparrow$) 4.24 3.16 3.04 3.04
Visual Quality($\uparrow$) 3.92 3.20 3.36 3.08
Table 3: The results of user study.
Ours SGSST ABC-GS StyleGaussian ARF Ref-NPR SNeRF
Memory(GB)($\downarrow$) 1.1 1.7 1.2 6.1 1.4 1.4 1.4
FPS($\uparrow$) 151 132 133 143 8 7 7
Table 4: Runtime Cost Comparison on the LLFF Dataset.
Figure 7: Ablation Study of Content Loss.
Figure 8: Qualitative Results of Texture Orientation Control.
Figure 9: Comparison of Qualitative Results.
Figure 10: Comparison of Qualitative Results.
DISCUSSION AND LIMITATION
Our method primarily consists of three key components: Geometry-aware Texture Transfer Loss (GT2 Loss), Adaptive Fine-grained Control Module (AFCM), and Geometry Preservation Branch (GPB). Among them, the proposed GT2 Loss and AFCM can be seamlessly applied to other representations, such as NeRF, TensoRF, and Plenoxel. Meanwhile, GPB provides a novel perspective for geometry-aware optimization in 3DGS-based appearance editing tasks without ground truth supervision. However, from another perspective, the upper bound of geometric accuracy after texture transfer in our framework depends on the quality of geometry obtained from the original 3DGS optimization, which may be suboptimal.
FUTURE WORK
In future work, we plan to focus on two main aspects. First, we aim to enhance the geometry correction stage of GPB by integrating more accurate 3DGS optimization techniques for geometry reconstruction. This improves the quality of scene appearance editing. In addition, we will explore the integration of diffusion models and multimodal large language models (MLLM) with 3D texture and style transfer. Leveraging the rich prior knowledge of multimodal models, we envision a unified framework that supports both text-driven and image-driven style transfer in 3D scenes.

### v4

@@ Introduction
3D style transfer (Zhang et al. 2022, 2024) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information. With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has increased significantly. 3D stylization techniques offer a promising solution by accelerating the creation of 3D content, particularly in complex 3D scene environments.
Recently, the emergence of Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) has significantly advanced the field of 3D stylization. NeRF-based stylization methods (Zhang et al. 2022, 2023; Nguyen-Phuoc, Liu, and Xiao 2022), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry. In contrast, 3DGS-based approaches (Galerne et al. 2025; Liu et al. 2025, 2024) offer benefits such as explicit editability and real-time rendering. 3D stylization methods can be broadly categorized into feed-forward and optimization-based approaches, depending on whether they support zero-shot style transfer. While feed-forward methods (Liu et al. 2023; Huang et al. 2022; Liu et al. 2024) enable zero-shot stylization, they often exhibit lower rendering quality. In contrast, optimization-based methods (Zhang et al. 2022; Liu et al. 2025; Zhang et al. 2024; Galerne et al. 2025) typically produce higher-fidelity stylization results. Despite these differences, both types of methods commonly define optimization objectives within the VGG feature space during training.
@@
Early texture transfer algorithms (Ashikhmin 2003; Lee et al. 2010; Hertzmann et al. 2023; Efros and Freeman 2023) were based on non-parametric texture synthesis. These methods preserve the structural information of the target image to achieve texture transfer effects. Gatys et al. (Gatys, Ecker, and Bethge 2016) pioneered the use of convolutional neural networks (CNNs) (Simonyan and Zisserman 2014) for image style transfer, achieving impressive results. Subsequently, a variety of deep learning-based style transfer methods (Chen and Schmidt 2016; Li and Wand 2016; Huang and Belongie 2017; Huo et al. 2021) have been proposed. These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer. Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (Vaswani et al. 2017), some deep learning-based texture transfer methods (Wang et al. 2022; Chen, Yin, and Fidler 2022; Pu et al. 2024) can transfer local textures from the reference image to the target image based on high-level semantic correspondence. In this work, we explore how to directly transfer texture features into 3D representations.
@@
3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency. Early research in 3D scene style transfer focused on explicit representations such as point clouds (Cao et al. 2020; Huang et al. 2021), meshes (Höllein, Johnson, and Nießner 2022; Yin et al. 2021), and voxels (Guo et al. 2021; Klehm et al. 2014). Zhang et al. (Zhang et al. 2022) explored the potential of NeRF in 3D stylization tasks. By optimizing the scene representation with the proposed 3D style transfer loss, it achieved pleasing visual effects. Subsequently, a variety of NeRF-based 3D style transfer methods (Huang et al. 2022; Pang, Hua, and Yeung 2023; Liu et al. 2023; Zhang, Fernandez-Labrador, and Schroers 2024; Jung et al. 2024) have emerged. Recently, the emergence of 3DGS (Kerbl et al. 2023) has brought new possibilities for scene stylization. It features fast training speed, high rendering quality, and efficient performance. Some works (Liu et al. 2024; Zhang et al. 2024; Mei, Xu, and Patel 2024; Liu et al. 2025) have already explored the potential of stylization under this representation. Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (Liu et al. 2024) and optimization-based approaches (Zhang et al. 2024; Mei, Xu, and Patel 2024; Liu et al. 2025). Feed-forward methods enable zero-shot style transfer. For instance, StyleGaussian (Liu et al. 2024) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (Huang and Belongie 2017) to achieve real-time stylization. On the other hand, optimization-based methods attain higher-quality style transfer results. Some works (Zhang et al. 2024; Liu et al. 2025) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization. However, existing methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images. Meanwhile, these methods overlook the fine-grained discrepancies between features and pixels. By incorporating geometric information and adaptive fine-grained control, our method is able to achieve high-quality texture transfer results.
@@
3D Gaussian Splatting (Kerbl et al. 2023) is a novel explicit 3D representation method. It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes. The parameters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, an opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$, which is represented as spherical harmonic coefficients. Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process. The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$. The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (Lassner and Zollhofer 2021). Specifically, the Gaussians are grouped into tiles and sorted by depth. The color $C$ of a pixel is computed through $\alpha$-blending:
@@
By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized. Existing optimization-based 3D style transfer methods (Zhang et al. 2022; Pang, Hua, and Yeung 2023; Zhang, Fernandez-Labrador, and Schroers 2024; Zhang et al. 2024) typically adopt the nearest neighbor feature matching (NNFM) loss (Zhang et al. 2022) as their style loss function. It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them. Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$. The same feature extractor (e.g., VGG (Simonyan and Zisserman 2014)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively. Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
@@
Datasets. For the scene datasets, we utilize the LLFF dataset (Mildenhall et al. 2019) and the Tanks and Temples (T&T) dataset (Knapitsch et al. 2017), which are collected from the real world. Additionally, we use images from the ARF (Zhang et al. 2022) style dataset and the DTD dataset (Cimpoi et al. 2014) as reference image datasets.
Baseline. We compare our method with the state-of-the-art 3D stylization methods, including SGSST (Galerne et al. 2025), ABC-GS (Liu et al. 2025), StyleGaussian (Liu et al. 2024), ARF (Zhang et al. 2022), Ref-NPR (Zhang et al. 2023), and SNeRF (Nguyen-Phuoc, Liu, and Xiao 2022). Specifically, SGSST, ABC-GS, and StyleGaussian are based on 3DGS, and ARF, Ref-NPR, and SNeRF are based on NeRF. StyleGaussian is a feed-forward-based method, while others are optimization-based methods.
@@
We perform view-consistent color transfer (Zhang et al. 2022) on the rendered images before and after texture transfer and use them as content images to optimize the Gaussian parameters. For the VGG (Simonyan and Zisserman 2014) feature extractor, we employ the conv3 block of VGG-16. When associating color parameters with texture features, we set the default number of depth groups $K$ to 4, and the rotation angle $\theta$ is sampled across the full 360 degrees. The hyperparameter for AFCM is denoted as $\{\lambda_{d},\lambda_{f},\lambda_{\Phi}\}$ = {0.8, 0.8, 0.25}. For texture transfer optimization, we set $\{\lambda_{wgt},\lambda_{c},\lambda_{tv}\}$ = {2, 0.005, 0.02}. All our experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@
Multi-view Consistency. We evaluate multi-view consistency (Liu et al. 2024) using both short-term and long-term consistency metrics. The results are shown in Tab. 1. It can be observed that both our method and ABC-GS achieve high-quality multi-view consistency. Notably, ABC-GS disables the Gaussian densification strategy during optimization. In contrast, our method maintains multi-view consistency even after applying the densification strategy, demonstrating the effectiveness of the proposed GPB.
Content Preservation. For 3D texture transfer, it is essential to ensure that the original scene content remains recognizable while editing the scene’s appearance. We use SSIM (Wang et al. 2004) and CLIP-score (Radford et al. 2021) to assess the preservation of content information. Specifically, SSIM evaluates the structural and informational similarity between two images at the pixel level. In contrast, CLIP-score measures the semantic similarity by computing the cosine similarity between the CLIP embeddings of the two images. As shown in Tab. 1, our proposed method achieves significantly higher scores on both evaluation metrics. In particular, our method outperforms previous approaches in terms of the CLIP-score. It indicates that incorporating geometric information enables accurate texture transfer to the scene appearance while preserving the semantic content of the scene.